The specific heat of Zn is 0.096 cal./g C. Determine the energy required to raise the temperature of 250 g of ZN from 20 degree C to 150 degrees C
1
Expert's answer
2018-01-24T12:15:07-0500
c = 0.096 cal./g C
m = 250 g
t1 = 20 C
t2 = 150 C
Q = c*m*(t2 - t1) = 0.096 cal./g C*250 g*(150 C -20 C) = 3120 cal.